# Submodular Maximization over a Matroid $k$-Intersection: Multiplicative Improvement over Greedy

### v1

@@ Abstract
All of our results hold also when the $k$-matroid intersection constraint is replaced with a more general matroid $k$-parity constraint. Furthermore, unlike the case in many of the previous works, our algorithms run in time that is independent of $k$ and polynomial in the size of the ground set. Our algorithms are based on a hybrid greedy local search approach recently introduced by Singer and Thiery [38] for the weighted matroid $k$-intersection problem, which is a special case of the problem we consider. Leveraging their approach in the submodular setting requires several non-trivial insights and algorithmic modifications since the marginals of a submodular function $f$, which correspond to the weights in the weighted case, are not independent of the algorithm’s internal randomness. In the special weighted case studied by [38], our algorithms reduce to a variant of the algorithm of [38] with an improved approximation ratio of $k\ln 2+1-\ln 2+\varepsilon<0.694k+0.307+\varepsilon$, compared to an approximation ratio of $\frac{k+1}{2\ln 2}\approx 0.722k+0.722$ guaranteed by Singer and Thiery [38].
@@ 1 Introduction
Improving over the above guarantees of the greedy algorithm has been an open question for many years, and until recently success has mostly been limited to special cases such as $k$-dimensional matching constraints. Here, we are given a hypergraph $H=(V,E)$ such that the set $V$ of vertices can be partitioned into $k$ disjoint subsets (or parts) $V_{1},V_{2}\dotsc,V_{k}$, and each hyperedge $e\in E$ contains at most 11 vertex from each one of the subsets $V_{i}$. Such hypergraphs are known as $k$-partite hypergraphs. The $k$-dimensional matching constraint requires that we select a set of vertex disjoint hyperedges from $E$ (i.e., a matching in $H$). This constraint can be captured by an intersection of $k$ partition matroid constraints, each enforcing that no vertex of a particular set $V_{i}$ is contained in two hyperedges.
@@
For general matroid $k$-intersection constraints, Lee, Sviridenko and Vondrák [28] gave a local search algorithm with an approximation ratio of k1k-1 for linear functions and $k$ for monotone submodular functions. In the special case of a cardinality objective, the same authors [29] gave a different local search algorithm with an approximation ratio of $k/2+\varepsilon$. We note that while the improvement in approximation for the cardinality objective has been by a multiplicative factor of 22, the improvements for submodular and linear functions have, until recently, been only by an additive term. This gap stems from the fact that the analysis presented in [29] is based on global arguments that are appropriate for the unweighted case in which all elements are equally valuable, but obtaining a result for a weighted problem requires more local arguments.
@@ 1.1 Our Results
Previous works [28, 42, 6, 8] have adapted local search algorithms for weighted maximization to submodular objectives by using the marginal or incremental contributions of an element $e$ as a surrogate weight. In other words, if $f$ is the objective function, then the adaptations implicitly define the weight of $e$ to be $w(e)\triangleq f(A\cup\{e\})-f(A)$, where AA is the current solution. [^2] In some cases, AA is in fact some appropriate subset of the current solution, but this distinction is unimportant for the sake of this high-level discussion. However, applying this idea to the algorithm of Singer and Thiery [38] is problematic.
@@
A significant bottleneck in the analysis of Singer and Thiery [38] occurs when there is a need to charge elements of the optimal solution to elements of the output solution that have nearly equal weight, but nevertheless belong to different weight classes. In this case, Singer and Thiery’s approximation guarantee degenerates to $k$, matching the guarantee of the greedy algorithm. This bottleneck affects only optimal elements whose weight lies at the end of the range of possible weights for their weight class. Thus, to overcome it Singer and Thiery apply a random shift to all weight classes, which allows them to bound the probability that any given optimal element has this undesirable property. In our setting, the surrogate “weight” of an element depends on the current solution AA being constructed by the algorithm, and thus is not independent of the random shift chosen; which means that it is no longer possible to bound the probability that an optimal element has a weight lying close to the end of the range of its weight class.
@@
Our solution for the above hurdle is to introduce an additional auxiliary weight $u(e)$ for every element $e$ in the optimal solution $O$. The weights $u(e)$ depend only on $O$, rather than on the solution set AA, and thus are independent of the random bits of the algorithm. However, as these weights depend on the unknown set $O$, they are unavailable to the algorithm, which must continue to use the weights $w(e)$. This limitation of the algorithm is not important when $w(e)$ and $u(e)$ roughly agree, and thus, the output of the algorithm is good (in expectation) in this case. In order to obtain a result that holds in general, we further show that whenever the two weights disagree significantly, this discrepancy can be used to obtain an alternative bound on the value of the output of the algorithm.
@@
Along the way, we also slightly modify the distribution of the random shift used by the algorithm to partition the elements into weight classes, which leads to some improvement also for linear objective functions compared to the approximation ratio of roughly $0.722k+O(1)$ obtained by [38]. More formally, we prove the following theorem. {restatable}theoremthmMonotoneResult For every $\varepsilon>0$, there exists an algorithm that runs in $\operatorname{Poly}(|E|,\varepsilon^{-1})$ time and guarantees an approximation ratio of $\frac{2k\ln 2}{1+\ln 2}+O(\sqrt{k})\leq 0.819k+O(\sqrt{k})$ for maximizing a non-negative monotone submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$ under a matroid $k$-parity constraint. Moreover, if $f$ is linear, the approximation ratio of the algorithm improves to $k\ln 2+1-\ln 2+\varepsilon\leq 0.694k+0.307+\varepsilon$.
@@
We also consider the maximization of a non-monotone submodular function subject to a matroid $k$-parity constraint. Algorithms for maximizing monotone submodular functions usually still produce for a non-monotone submodular function $f$ a guarantee of the form $f(A)\geq\alpha\cdot f(A\cup O)$, where AA is the output set of the algorithm, $O$ is the optimal solution and $\alpha$ is positive number. Gupta, Roth, Schoenebeck and Talwar [18] described a general framework for translating such a guarantee into a real approximation ratio, and their framework was improved by subsequent works [33, 13]. Unfortunately, because of the use of our novel auxiliary weights, when our algorithm from Theorem 1.1 is analyzed in the context of a non-monotone submodular function $f$, it yields a guarantee of the more involved form $f(A)\geq\alpha\cdot f(A\cup O)-\beta\cdot f(O)$ for some $\beta<\alpha$. Nevertheless, we are able to adapt the version of the framework used by [13] to transform guarantees of our more involved form into standard approximation ratios, which yields the following theorem. {restatable}theoremthmNonMonotoneResult There exists an algorithm that runs in $\operatorname{Poly}(|E|)$ time and guarantees an approximation ratio of $\frac{2k\ln 2}{1+\ln 2}+O(k^{2/3})\leq 0.819k+O(k^{2/3})$ for maximizing a non-negative (not necessarily monotone) submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$ under a matroid $k$-parity constraint.
@@ 1.2 Additional Related Work
When $k=2$, one can exactly optimize a linear function over a matroid $k$-intersection constraint via the classical algorithms of Edmonds [12] or Cunningham [10]. Similarly, a 22-set packing constraint is simply a standard matching constraint, and one can exactly optimize linear functions over it in polynomial time. Lawler [26] introduced the family of matroid parity constraints as a common generalization of both these families of constraints. Here, we are given a set $E$ of disjoint pairs of elements from $V$ and a matroid $\mathcal{M}$} on $V$, and the goal is to find a collection of pairs from $E$ whose union is independent for $\mathcal{M}$}. In contrast to the constraints they generalize, it is impossible to maximize even a cardinality objective under a matroid parity constraint in polynomial time in the setting in which the matroid is given as an independence oracle [31, 23]. Maximizing a cardinality objective under a matroid parity constraint is also known to be NP-hard (see, e.g., [37, §43.9]) and Lee, Sviridenko and Vondrák [29] showed that it admits a PTAS.
When $k\geq 3$, optimizing over both matroid $k$-intersection and $k$-set packing constraints is NP-hard as both these constraint families generalize $k$-dimensional matching constraints, and optimizing over the last family of constraints is NP-hard even for cardinality objectives [24]. All these types of constraints are generalized by the family of matroid $k$-parity constraints, and this family also generalizes the above-mentioned family of matroid parity constraints in the sense that matroid parity constraints are equivalent to matroid 22-parity constraints. The formal definition of matroid $k$-parity constraints can be found in Section 2, and we refer the reader to [29] for a more detailed discussion of the relationship between these and other families of constraints.
Finally, we note that improved results are known in several special cases. For matroid 22-intersection constraints, Huang and Sellier [20] give a $\frac{3}{2}$-approximation in the special case that the objective is a vertex cover function. For matroid $k$-parity constraints, there are several results for the special case that the matroid is linear and a representation is given. In this setting, when $k=2$ an exact polynomial time algorithm for cardinality objectives was given by Lovász [30], and Iwata and Kobayashi [21] later gave an exact polynomial time algorithm for maximizing linear objectives. For larger $k$ values, in the same setting, exact FPT algorithms are known for both cardinality and linear objectives, with exponential dependence on $k$ and the rank of the matroid or the number of blocks in the solution [2, 32]. In contrast, in this paper, we give approximation algorithms that require no assumptions on the structure of the underlying matroid and have a running time that is independent of $k$.
@@ 2.1 Constraint Preliminaries
Let ${\mathcal{M}}=(V,{\mathcal{I}}_{\mathcal{M}})$ and $E\subseteq 2^{V}$ be the matroid and partition defining the given matroid $k$-parity constraint. Suppose first that $A\cap B\neq\emptyset$. Then, we can define a new matroid $k$-parity constraint based on the matroid ${\mathcal{M}}/v(A\cap B)$ and the set of edges $E\setminus A\cap B$. Clearly, $A\setminus B$ and $B\setminus A$ are disjoint and feasible for this new matroid $k$-parity constraint. By applying the present lemma to these disjoint sets and this new constraint, we can obtain a collection of sets $\{N_{b}\subseteq A\setminus B\mid b\in B\setminus A\}$. Furthermore, adding to this collection the set $N_{b}=\{b\}$ for every $b\in A\cap B$ yields an extended collection of sets obeying all the properties required by the present lemma for the original sets AA and $B$ and the original matroid $k$-parity constraint. Thus, to prove the present lemma for general sets AA and $B$, it suffices to prove it for disjoint sets AA and $B$, which is our goal in the rest of this proof.
@@
Claim 1 follows immediately from our assumption that AA and $B$ are disjoint. Let us prove next claim 2. Define $S=\{b\in B\mid N_{b}=\emptyset\}$. For every $b\in S$, it holds by definition that $\pi_{a}\cap\overline{v}(b)=\emptyset$ for every $a\in A$. Since the sets $\pi_{a}$ are a partition of $\overline{v}(B)$ and $\overline{v}(b)\subseteq\overline{v}(B)$, we must have $\overline{v}(b)=\emptyset$ for every such $b\in S$. Thus,
@@
For the proof of claim 3, fix $a\in A$ and let $S=\{b\in B\mid N_{b}=\{a\}\}$. For all $b\in S$, $\overline{v}(b)\subseteq\pi_{a}$ because otherwise $\overline{v}(b)$ would have intersected additional sets in the partition of $\overline{v}(B)$ beside $\pi_{a}$, which would have resulted in additional elements of AA appearing in $N_{b}$. Thus,
@@ 3 Main Algorithm
In this section, we present our main algorithm, given below as Algorithm 1. In Section 4, we prove that Algorithm 1 provides the approximation guarantees for linear and monotone submodular functions given in Theorem $1.$ Algorithm 1 is also a central component in the proof of Theorem 1.1 for non-monotone submodular objectives. As stated, Algorithm 1 is not necessarily a polynomial time algorithm. However, in Appendix A we describe a polynomial time implementation of it.
@@
Let $\theta>0$ and $\varepsilon\in(0,1)$ be two given values, and $A\in{\mathcal{I}}$ be a feasible solution. We say that a pair of sets $S\subseteq E\setminus A$ and $N\subseteq A$ is a $(\theta,\varepsilon)$-improvement for AA if $(A\cup S)\setminus N\in{\mathcal{I}}$ and one of the following holds:
@@
At the beginning of iteration $i$ of Algorithm 1, the sets $A_{1},A_{2},\dotsc,A_{i-1}$ are already finalized. Algorithm 1 then tries to construct a set $A_{i}$ that is a good addition to these sets. It does so by initializing $A_{i}$ to be the empty set, and then improving $A_{\leq i-1}\cup A_{i}$ via a series of $(m_{i},\varepsilon)$-improvements. Notice that, by definition, the elements added by these improvements to $A_{i}$ have marginals of at least $m_{i}$, as promised above. Since the sets $A_{1},A_{2},\dotsc,A_{i-1}$ are finalized before the $i$-th iteration of Algorithm 1, and must not be modified at this iteration, the algorithm considers only $m_{i}$-improvements whose set NN of removed elements is a subset of the current set $A_{i}$. Whenever an improvement of this kind is found, it is applied by setting $A_{i}\leftarrow(A_{i}\cup S)\setminus N$, and once no more such $m_{i}$-improvements can be found, the $i$-th iteration of Algorithm 1 terminates.
@@
In Appendix A, we show that Algorithm 1 can in fact be simulated in polynomial time. However, in order to derive the approximation guarantees of Theorem 1.1, it is sufficient to prove here the weaker claim that Algorithm 1 terminates. In each iteration $i$, Algorithm 1 repeatedly searches for and applies $(m_{i},\varepsilon)$-improvements for some current solution AA. Each such improvement either increases $|A|$, or leaves $|A|$ unchanged and increases $f(A)$. Since there are only finitely many possible distinct values for $|A|$ and $f(A)$, it follows that each iteration of Algorithm 1 must terminate. The following observation implies that Algorithm 1 can perform at most finitely many iterations before terminating.
@@
First, we note that any iteration $i$ in which at least one $(m_{i},\varepsilon)$-improvement is found, must end with $|A_{i}|>0$. This follows directly from the structure of $(\theta,\varepsilon)$-improvements, which never decrease the size of the set AA that is being improved. Since the sets $A_{i}$ produced across different iterations are disjoint by construction, there can be at most $|E|$ iterations in which the algorithm finds at least one $(\theta,\varepsilon)$-improvement. Let $L$ be the index of the last such iteration, or 0 if there are no such iterations. Suppose, for the sake of contradiction, that there is some $e\in E\setminus A_{\leq L}$ with $A_{\leq L}+e\in{\mathcal{I}}$ and $f(e\mid A_{\leq L})>0$. Since no $(\theta,\varepsilon)$-improvements are found after iteration $L$, we must have $A_{\leq i}=A_{\leq L}$ for all $i>L$. But, for sufficiently large $i\geq 1+\log_{2}W-\log_{2}f(e\mid A_{\leq L})$, we then have $e\in E\setminus A_{\leq i}$ with $f(e\mid A_{\leq i})=f(e\mid A_{\leq L})\geq m_{i}$ and $A_{\leq i}+e=A_{\leq L}+e\in{\mathcal{I}}$. Thus, $(\{e\},\emptyset)$ would be a valid $(m_{i},\varepsilon)$-improvement in iteration $i>L$, contradicting the fact that no such improvements are found by Algorithm 1. ∎
@@
For every $i\in[L]$, the inequality $f(e\mid A_{\leq i})<m_{i}$ holds for every $e\in E\setminus A_{\leq i}$ that obeys $A_{\leq i}+e\in{\mathcal{I}}$ at the end of iteration ii.
@@
Suppose $A_{\leq i}+e\in{\mathcal{I}}$ for some $e\in E\setminus A_{\leq i}$ at the end of iteration ii. If $f(e\mid A_{\leq i})\geq m_{i}$, then $(\{e\},\emptyset)$ would be a valid $(m_{i},\varepsilon)$-improvement, contradicting the fact that the loop on Line 1 of Algorithm 1 terminated. ∎

### v2

@@ Abstract
All of our results hold also when the $k$-matroid intersection constraint is replaced with a more general matroid $k$-parity constraint. Furthermore, unlike the case in many of the previous works, our algorithms run in time that is independent of $k$ and polynomial in the size of the ground set. Our algorithms are based on a hybrid greedy local search approach recently introduced by Singer and Thiery [38] for the weighted matroid $k$-intersection problem, which is a special case of the problem we consider. Leveraging their approach in the submodular setting requires several non-trivial insights and algorithmic modifications since the marginals of a submodular function $f$, which correspond to the weights in the weighted case, are not independent of the algorithm’s internal randomness. In the special weighted case studied by [38], our algorithms reduce to a variant of the algorithm of [38] with an improved approximation ratio of $(k+1)\ln 2+O(\varepsilon)<0.694k+0.694+O(\varepsilon)$, compared to an approximation ratio of $\frac{k+1}{2\ln 2}\approx 0.722k+0.722$ guaranteed by Singer and Thiery [38].
@@ 1 Introduction
Improving over the above guarantees of the greedy algorithm has been an open question for many years, and until recently success has mostly been limited to special cases such as $k$-dimensional matching constraints. Here, we are given a hypergraph $H=(V,E)$ such that the set $V$ of vertices can be partitioned into $k$ disjoint subsets (or parts) $V_{1},V_{2}\dotsc,V_{k}$, and each hyperedge $e\in E$ contains at most $1$ vertex from each one of the subsets $V_{i}$. Such hypergraphs are known as $k$-partite hypergraphs. The $k$-dimensional matching constraint requires that we select a set of vertex disjoint hyperedges from $E$ (i.e., a matching in $H$). This constraint can be captured by an intersection of $k$ partition matroid constraints, each enforcing that no vertex of a particular set $V_{i}$ is contained in two hyperedges.
@@
For general matroid $k$-intersection constraints, Lee, Sviridenko and Vondrák [28] gave a local search algorithm with an approximation ratio of k1k-1 for linear functions and $k$ for monotone submodular functions. In the special case of a cardinality objective, the same authors [29] gave a different local search algorithm with an approximation ratio of $k/2+\varepsilon$. We note that while the improvement in approximation for the cardinality objective has been by a multiplicative factor of $2$, the improvements for submodular and linear functions have, until recently, been only by an additive term. This gap stems from the fact that the analysis presented in [29] is based on global arguments that are appropriate for the unweighted case in which all elements are equally valuable, but obtaining a result for a weighted problem requires more local arguments.
@@ 1.1 Our Results
Previous works [28, 42, 6, 8] have adapted local search algorithms for weighted maximization to submodular objectives by using the marginal or incremental contributions of an element $e$ as a surrogate weight. In other words, if $f$ is the objective function, then the adaptations implicitly define the weight of $e$ to be $w(e)\triangleq f(A\cup\{e\})-f(A)$, where $A$ is the current solution. [^2] In some cases, $A$ is in fact some appropriate subset of the current solution, but this distinction is unimportant for the sake of this high-level discussion. However, applying this idea to the algorithm of Singer and Thiery [38] is problematic.
@@
A significant bottleneck in the analysis of Singer and Thiery [38] occurs when there is a need to charge elements of the optimal solution to elements of the output solution that have nearly equal weight, but nevertheless belong to different weight classes. In this case, Singer and Thiery’s approximation guarantee degenerates to $k$, matching the guarantee of the greedy algorithm. This bottleneck affects only optimal elements whose weight lies at the end of the range of possible weights for their weight class. Thus, to overcome it Singer and Thiery apply a random shift to all weight classes, which allows them to bound the probability that any given optimal element has this undesirable property. In our setting, the surrogate “weight” of an element depends on the current solution $A$ being constructed by the algorithm, and thus is not independent of the random shift chosen; which means that it is no longer possible to bound the probability that an optimal element has a weight lying close to the end of the range of its weight class.
@@
Our solution for the above hurdle is to introduce an additional auxiliary weight $u(e)$ for every element $e$ in the optimal solution $O$. The weights $u(e)$ depend only on $O$, rather than on the solution set $A$, and thus are independent of the random bits of the algorithm. However, as these weights depend on the unknown set $O$, they are unavailable to the algorithm, which must continue to use the weights $w(e)$. This limitation of the algorithm is not important when $w(e)$ and $u(e)$ roughly agree, and thus, the output of the algorithm is good (in expectation) in this case. In order to obtain a result that holds in general, we further show that whenever the two weights disagree significantly, this discrepancy can be used to obtain an alternative bound on the value of the output of the algorithm.
@@
Along the way, we also slightly modify the distribution of the random shift used by the algorithm to partition the elements into weight classes, which leads to some improvement also for linear objective functions compared to the approximation ratio of roughly $0.722k+O(1)$ obtained by [38]. More formally, we prove the following theorem. {restatable}theoremthmMonotoneResult For every $\varepsilon>0$, there exists an algorithm that runs in $\operatorname{Poly}(|E|,\varepsilon^{-1})$ time and guarantees an approximation ratio of $\frac{2k\ln 2}{1+\ln 2}+O(\sqrt{k})\leq 0.819k+O(\sqrt{k})$ for maximizing a non-negative monotone submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$ under a matroid $k$-parity constraint. Moreover, if $f$ is linear, the approximation ratio of the algorithm improves to $(k+1)\ln 2+O(\varepsilon)\leq 0.694k+0.694+O(\varepsilon)$.
@@
We also consider the maximization of a non-monotone submodular function subject to a matroid $k$-parity constraint. Algorithms for maximizing monotone submodular functions usually still produce for a non-monotone submodular function $f$ a guarantee of the form $f(A)\geq\alpha\cdot f(A\cup O)$, where $A$ is the output set of the algorithm, $O$ is the optimal solution and $\alpha$ is positive number. Gupta, Roth, Schoenebeck and Talwar [18] described a general framework for translating such a guarantee into a real approximation ratio, and their framework was improved by subsequent works [33, 13]. Unfortunately, because of the use of our novel auxiliary weights, when our algorithm from Theorem 1.1 is analyzed in the context of a non-monotone submodular function $f$, it yields a guarantee of the more involved form $f(A)\geq\alpha\cdot f(A\cup O)-\beta\cdot f(O)$ for some $\beta<\alpha$. Nevertheless, we are able to adapt the version of the framework used by [13] to transform guarantees of our more involved form into standard approximation ratios, which yields the following theorem. {restatable}theoremthmNonMonotoneResult There exists an algorithm that runs in $\operatorname{Poly}(|E|)$ time and guarantees an approximation ratio of $\frac{2k\ln 2}{1+\ln 2}+O(k^{2/3})\leq 0.819k+O(k^{2/3})$ for maximizing a non-negative (not necessarily monotone) submodular function $f\colon 2^{E}\to{{\mathbb{R}}_{\geq 0}}$ under a matroid $k$-parity constraint.
@@ 1.2 Additional Related Work
When $k=2$, one can exactly optimize a linear function over a matroid $k$-intersection constraint via the classical algorithms of Edmonds [12] or Cunningham [10]. Similarly, a $2$-set packing constraint is simply a standard matching constraint, and one can exactly optimize linear functions over it in polynomial time. Lawler [26] introduced the family of matroid parity constraints as a common generalization of both these families of constraints. Here, we are given a set $E$ of disjoint pairs of elements from $V$ and a matroid $\mathcal{M}$} on $V$, and the goal is to find a collection of pairs from $E$ whose union is independent for $\mathcal{M}$}. In contrast to the constraints they generalize, it is impossible to maximize even a cardinality objective under a matroid parity constraint in polynomial time in the setting in which the matroid is given as an independence oracle [31, 23]. Maximizing a cardinality objective under a matroid parity constraint is also known to be NP-hard (see, e.g., [37, §43.9]) and Lee, Sviridenko and Vondrák [29] showed that it admits a PTAS.
When $k\geq 3$, optimizing over both matroid $k$-intersection and $k$-set packing constraints is NP-hard as both these constraint families generalize $k$-dimensional matching constraints, and optimizing over the last family of constraints is NP-hard even for cardinality objectives [24]. All these types of constraints are generalized by the family of matroid $k$-parity constraints, and this family also generalizes the above-mentioned family of matroid parity constraints in the sense that matroid parity constraints are equivalent to matroid $2$-parity constraints. The formal definition of matroid $k$-parity constraints can be found in Section 2, and we refer the reader to [29] for a more detailed discussion of the relationship between these and other families of constraints.
Finally, we note that improved results are known in several special cases. For matroid $2$-intersection constraints, Huang and Sellier [20] give a $\frac{3}{2}$-approximation in the special case that the objective is a vertex cover function. For matroid $k$-parity constraints, there are several results for the special case that the matroid is linear and a representation is given. In this setting, when $k=2$ an exact polynomial time algorithm for cardinality objectives was given by Lovász [30], and Iwata and Kobayashi [21] later gave an exact polynomial time algorithm for maximizing linear objectives. For larger $k$ values, in the same setting, exact FPT algorithms are known for both cardinality and linear objectives, with exponential dependence on $k$ and the rank of the matroid or the number of blocks in the solution [2, 32]. In contrast, in this paper, we give approximation algorithms that require no assumptions on the structure of the underlying matroid and have a running time that is independent of $k$.
@@ 2.1 Constraint Preliminaries
Let ${\mathcal{M}}=(V,{\mathcal{I}}_{\mathcal{M}})$ and $E\subseteq 2^{V}$ be the matroid and partition defining the given matroid $k$-parity constraint. Suppose first that $A\cap B\neq\emptyset$. Then, we can define a new matroid $k$-parity constraint based on the matroid ${\mathcal{M}}/v(A\cap B)$ and the set of edges $E\setminus A\cap B$. Clearly, $A\setminus B$ and $B\setminus A$ are disjoint and feasible for this new matroid $k$-parity constraint. By applying the present lemma to these disjoint sets and this new constraint, we can obtain a collection of sets $\{N_{b}\subseteq A\setminus B\mid b\in B\setminus A\}$. Furthermore, adding to this collection the set $N_{b}=\{b\}$ for every $b\in A\cap B$ yields an extended collection of sets obeying all the properties required by the present lemma for the original sets $A$ and $B$ and the original matroid $k$-parity constraint. Thus, to prove the present lemma for general sets $A$ and $B$, it suffices to prove it for disjoint sets $A$ and $B$, which is our goal in the rest of this proof.
@@
Claim 1 follows immediately from our assumption that $A$ and $B$ are disjoint. Let us prove next claim 2. Define $S=\{b\in B\mid N_{b}=\emptyset\}$. For every $b\in S$, it holds by definition that $\pi_{a}\cap\overline{v}(b)=\emptyset$ for every $a\in A$. Since the sets $\pi_{a}$ are a partition of $\overline{v}(B)$ and $\overline{v}(b)\subseteq\overline{v}(B)$, we must have $\overline{v}(b)=\emptyset$ for every such $b\in S$. Thus,
@@
For the proof of claim 3, fix $a\in A$ and let $S=\{b\in B\mid N_{b}=\{a\}\}$. For all $b\in S$, $\overline{v}(b)\subseteq\pi_{a}$ because otherwise $\overline{v}(b)$ would have intersected additional sets in the partition of $\overline{v}(B)$ beside $\pi_{a}$, which would have resulted in additional elements of $A$ appearing in $N_{b}$. Thus,
@@ 3 Main Algorithm
In this section, we present our main algorithm, given below as Algorithm 1. In Section 4, we prove that Algorithm 1 provides the approximation guarantees for linear and monotone submodular functions given in Theorem 1.1. Algorithm 1 is also a central component in the proof of Theorem 1.1 for non-monotone submodular objectives. As stated, Algorithm 1 is not necessarily a polynomial time algorithm. However, in Appendix A we describe a polynomial time implementation of it.
@@
Let $\theta>0$ and $\varepsilon\in(0,1)$ be two given values, and $A\in{\mathcal{I}}$ be a feasible solution. We say that a pair of sets $S\subseteq E\setminus A$ and $N\subseteq A$ is a $(\theta,\varepsilon)$-improvement for $A$ if $(A\cup S)\setminus N\in{\mathcal{I}}$ and one of the following holds:
@@
At the beginning of iteration $i$ of Algorithm 1, the sets $A_{1},A_{2},\dotsc,A_{i-1}$ are already finalized. Algorithm 1 then tries to construct a set $A_{i}$ that is a good addition to these sets. It does so by initializing $A_{i}$ to be the empty set, and then improving $A_{\leq i-1}\cup A_{i}$ via a series of $(m_{i},\varepsilon)$-improvements. Notice that, by definition, the elements added by these improvements to $A_{i}$ have marginals of at least $m_{i}$, as promised above. Since the sets $A_{1},A_{2},\dotsc,A_{i-1}$ are finalized before the $i$-th iteration of Algorithm 1, and must not be modified at this iteration, the algorithm considers only $m_{i}$-improvements whose set $N$ of removed elements is a subset of the current set $A_{i}$. Whenever an improvement of this kind is found, it is applied by setting $A_{i}\leftarrow(A_{i}\cup S)\setminus N$, and once no more such $m_{i}$-improvements can be found, the $i$-th iteration of Algorithm 1 terminates.
@@
In Appendix A, we show that Algorithm 1 can in fact be simulated in polynomial time. However, in order to derive the approximation guarantees of Theorem 1.1, it is sufficient to prove here the weaker claim that Algorithm 1 terminates. In each iteration $i$, Algorithm 1 repeatedly searches for and applies $(m_{i},\varepsilon)$-improvements for some current solution $A$. Each such improvement either increases $|A|$, or leaves $|A|$ unchanged and increases $f(A)$. Since there are only finitely many possible distinct values for $|A|$ and $f(A)$, it follows that each iteration of Algorithm 1 must terminate. The following observation implies that Algorithm 1 can perform at most finitely many iterations before terminating.
@@
First, we note that any iteration $i$ in which at least one $(m_{i},\varepsilon)$-improvement is found, must end with $|A_{i}|>0$. This follows directly from the structure of $(\theta,\varepsilon)$-improvements, which never decrease the size of the set $A$ that is being improved. Since the sets $A_{i}$ produced across different iterations are disjoint by construction, there can be at most $|E|$ iterations in which the algorithm finds at least one $(\theta,\varepsilon)$-improvement. Let $L$ be the index of the last such iteration, or 0 if there are no such iterations. Suppose, for the sake of contradiction, that there is some $e\in E\setminus A_{\leq L}$ with $A_{\leq L}+e\in{\mathcal{I}}$ and $f(e\mid A_{\leq L})>0$. Since no $(\theta,\varepsilon)$-improvements are found after iteration $L$, we must have $A_{\leq i}=A_{\leq L}$ for all $i>L$. But, for sufficiently large $i\geq 1+\log_{2}W-\log_{2}f(e\mid A_{\leq L})$, we then have $e\in E\setminus A_{\leq i}$ with $f(e\mid A_{\leq i})=f(e\mid A_{\leq L})\geq m_{i}$ and $A_{\leq i}+e=A_{\leq L}+e\in{\mathcal{I}}$. Thus, $(\{e\},\emptyset)$ would be a valid $(m_{i},\varepsilon)$-improvement in iteration $i>L$, contradicting the fact that no such improvements are found by Algorithm 1. ∎
@@
For every $i\in[L]$, the inequality $f(e\mid A_{\leq i})<m_{i}$ holds for every $e\in E\setminus A_{\leq i}$ that obeys $A_{\leq i}+e\in{\mathcal{I}}$ at the end of iteration $i$.
@@
Suppose $A_{\leq i}+e\in{\mathcal{I}}$ for some $e\in E\setminus A_{\leq i}$ at the end of iteration $i$. If $f(e\mid A_{\leq i})\geq m_{i}$, then $(\{e\},\emptyset)$ would be a valid $(m_{i},\varepsilon)$-improvement, contradicting the fact that the loop on Line 1 of Algorithm 1 terminated. ∎